OpenStudy (anonymous):

If \(\large \left|\frac{z_1 + z_2}{z_1 - z_2}\right| = 1\), \(z_2 \neq 0\) and \(z_1 \neq z_2\). Then what do we know about \(\frac{z_1}{z_2}\).

OpenStudy (anonymous):

I think we can use Componendo-Dividendo or whatever it's called. \[\frac{z_1}{z_2} = \pm 1\]But \(z_1 \neq z_2\), so only one case when \(z_1 = -z_2\).

OpenStudy (anonymous):

Hmm \(\Large\frac{z_1}{z_2}\) is purely real.

OpenStudy (lgbasallote):

\(\large \color{blue}{\mathtt{o}} \color{forestgreen}{\mathsf{h}}\) \(\LARGE \color{red}{\mathbb{\LaTeX}} \color{orange}{\mathcal{!}}\)

OpenStudy (anonymous):

Lol Nice. LGB!

OpenStudy (anonymous):

And negative as 1 isn't possible. Only -1.

OpenStudy (anonymous):

I think I have it. Thanks! lol

OpenStudy (anonymous):

\[\huge \mathsf{\color{goldenrod}{\text{Nice,}}}\quad \mathrm{\color{royalblue}{\LaTeX}}.\]

OpenStudy (lgbasallote):

\(\huge \color{gold}{\textbf{you're}}\) \(\huge \color{gray}{\mathsf{good}}\)